Fermi National Laboratory
 

Operation Summary for September 1, 2010 - September 3, 2010

 

Wednesday Sept. 1

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (3.4E10) and the Antiproton Source (Pbar) stack (26.3E10), with the Linac and Booster operating normally, with the Main Injector (MI) off due to a vacuum problem, with the MiniBooNE experiment taking beam, with NuMI off due to MI vacuum problems, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations established store 8047 at 8:15 AM with an average initial luminosity (AIL) of 204E30. Operators prepared MI, Booster, MiniBooNE, MI-65, and the MINOS absorber for access. Vacuum technicians found and repaired a vacuum leak at MI-520 on a beam tube flange near IP521C; by the end of the day shift they had started the pump down. Linac personnel replaced the auxiliary filament control breaker for a RF station (LRF2), but the station developed another problem; Linac personnel continued to investigate. RF personnel repaired a Booster RF station (BRF15). No beam was delivered to the users during this shift.

During the evening shift, MI and Booster personnel came out of access at 5 PM. Operations established beam to MiniBooNE at 6:15 PM and to Pbar and NuMI by 6:45 PM; Operations increased beam intensities as the vacuum improved at MI-520.

Thursday Sept. 2

The midnight shift began with Operations monitoring store 8047 (51.12E30) and the stash (123.9E10), with Pbar stacking (18.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that the vacuum at MI-520 continued to improve. Operations terminated store 8047 at 6 AM. TeV personnel conducted proton orbit studies.

During the day shift, Operations put the TeV into shot setup at 8:30 AM and established store 8051 at 9:45 AM with an AIL of 183E30. Unfortunately, store 8051 aborted at noon due to vacuum problems at sector F39. Mechanical Support personnel accessed F-sector and located a rather large and suspicious ice ball near the turn around box. After investigating further they reported that the ice ball did not appear to be the most significant vacuum problem in the area. Efforts to locate and repair the vacuum leak continued throughout the shift. Linac personnel began to work on the LRF2 filament control breaker at 1:30 PM. Operations prepared Enclosure B, F sector, Transfer Hall, and D0 for access.

During the evening shift, the F-Sector and Transfer Hall accesses ended at 8:30 PM. Operations established beam to MiniBooNE at 8:45 PM and to NuMI shortly thereafter. Beam to the Pbar Source was delayed due to power supply (F17B3) regulation problems. The regulation problem resolved itself as EE Support personnel arrived to investigate. Pbar resumed stacking at 9:40 PM. A weather-related power glitch caused various RF systems to trip and a 72E10 loss of antiprotons in the Recycler.

Friday Sept. 3

The midnight shift began with the TeV down due to a vacuum leak in F sector, with Operations monitoring the stash (148.6E10), with Pbar stacking (17.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators prepared F-sector for access at 6:30 AM to continue with vacuum repairs; with F-sector in access there is no MI beam. Operations reported some minor downtime due to LRF5.

The Plans for Friday and the Weekend

The plans for today are to continue the TeV F-sector vacuum work and then resume stacking, stashing, and storing for today and the weekend. The Run Coordinator will allow MI personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory